Benzene (C6H6) and cyclohexane (C6H12) form an ideal solution at 25°C. The density of benzene is 0.876 g/cm 3 and the density of cyclohexane is unknown.A liquid mixture is prepared by mixing 357 mL of benzene and 643 mL of cyclohexane. The vapor in equilibrium with this liquid has a mole fraction of benzene of 0.4. What is the density of cyclohexane in units of g/cm3?

Data:

Vapor pressure of benzene = 99 mmHg, at 25°C

Vapor pressure of cyclohexane = 100 mmHg, at 25°C

1) 0.70

2) 0.77

3) 0.87

4) 0.92

5) 1.07
Answer #1

Ans:- Here mole fraction of benzene XB = 0.4

Therefore mole fraction of cyclohexane XA = 1 - 0.4=0.6

Given Vapour pressure of Cyclohexane is P0A = 100mmhg

Density of benzene = mass of benzene / volume

0.876                                     = mass of benzene WB / 357

WB = 31.27g

Molar mass of benzene MB = 78

Molar mass of cyclohexane = 84

Now , using raoults law for ideal solution

MB = WB x MA / WA x   XA

78 = 31.27 x 84 / WA x 0.6

WA =46.8

density of cyclohexane = 46.8/64.3

                                   = 0.87 g/cm3

The option (3) is correct